OpenStudy (anonymous):

Select next number in series: Please help!!! 60-30-20-15-12 a. 13 b. 6 c. 10 d. 4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

pattern I found : 60*1/2 = 30 30*2/3 = 20 20*3/4 = 15 15*4/5 = 12 12*5/6 = 10

OpenStudy (anonymous):

So your answer is 10? That is the same as I had, but a friend of mine said the answer was 4 & got me confused. Thank you & if you notice...mexvizu answered but I will stick with 10 which was my original answer.

OpenStudy (anonymous):

so, you were confused 'cause your friend said answer is 4 without explanation ...?